# intrinsic motivation; active inference; empowerment; entropy

### v2

@@ 1 Introduction
In psychology, “intrinsic motivation” refers to the tendency for intelligent creatures to be motivated to do certain things (such as explore, learn, and grow) even in the absence of specific external reward signals (Domenico and Ryan, 2017). Recently, this paradigm has gained traction in machine learning, where it is operationalized as the idea that policies for action may be optimized based on structural features of agents and agent-environment interactions, as against traditional approaches like reinforcement learning, which optimize policies based on ad hoc reward functions. [^1] For pioneering work on intrinsic motivation in artificial intelligence, see (Schmidhuber, 2010).
@@
The active inference framework (Friston et al., 2017) shares similar motivations, and provides a Bayesian framework for combining a general form of intrinsic motivation (i.e. curiosity or “epistemic drive”) (Friston et al., 2015) with agent-specific prior distributions over states or outcomes (Da Costa et al., 2024), which model homeostatic set points and can function like explicit rewards. The expected (variational) free energy (EFE), as discussed below, guides policy selection in this framework by supplying an empirical prior over policies, given observations.
@@
2 Formal accounts of intrinsic motivation

### v3

@@ 1 Introduction
In psychology, “intrinsic motivation” refers to the tendency for intelligent creatures to be motivated to do certain things (such as explore, learn, and grow) even in the absence of specific external reward signals (Domenico and Ryan, 2017). This paradigm has increasingly gained traction in machine learning, where it is operationalized as the idea that policies for action may be optimized based on structural features of agents and agent-environment interactions, as against traditional approaches like reinforcement learning, which optimize policies based on ad hoc reward functions.
@@
The active inference framework (Friston et al., 2017) shares similar motivations, and provides a Bayesian method for combining a general form of intrinsic motivation (i.e. curiosity or “epistemic drive”) (Friston et al., 2015) with agent-specific prior distributions over states or outcomes (Da Costa et al., 2024), which model homeostatic set points and can function like explicit rewards. The expected (variational) free energy (EFE), as discussed below, guides policy selection in this framework by supplying an empirical prior over policies, given observations.
@@
There are many other formal treatments of intrinsic motivation in the literature on machine learning, some closely related to those just discussed, such as pioneering work on artificial curiosity (see e.g. Schmidhuber (1991, 2010)) and treatments in terms of Bayesian surprise (Itti and Baldi, 2009; Mazzaglia et al., 2021). Here, the focus is mainly on the relationship between active inference and empowerment, and on the relationship of both to maximum occupancy which has recently been proposed explicitly as an alternative.
@@
2 Three formal accounts of intrinsic motivation
@@ 2.2 Active inference and expected free energy
Among the most promising approaches to intrinsic motivation are those that leverage Bayesian